paid P1,080 in interest on a P3,060 loan she got from the credit union. The term of the

loan is 3 years. Find the rate.


1
Expert's answer
2021-12-13T05:58:10-0500

"r=\frac{\frac{Si}{S}-1}{n}"


Si=1080+3060=4140

S=3060

n=3

"r=\frac{\frac{Si}{S}-1}{n}=\frac{\frac{4140}{3060}-1}{3}=0.117" or 11.7%